OpenStudy (anonymous):

Can someone please simplify, (with illustrative steps) the following expression: 2sqrt20x + 3sqrt5x? Thanks!

OpenStudy (larseighner):

It appears to me you take a factor of 4 out of the first radical.

OpenStudy (anonymous):

\[2\sqrt20x +3\sqrt{5x}\]

OpenStudy (mathstudent55):

\(2\sqrt{20x} +3\sqrt{5x}\) To simplify the left square root, find the largest perfect square factor of the number part. The number part there is 20. What is the largest perfect square factor of 20?

OpenStudy (anonymous):

ok so 4 and 5 right? Then what?

OpenStudy (larseighner):

You can take the 4 out of the radical. On the outside, of course, it becomes its square root. I don't think it is a spoiler to say the root 4 is 2,

OpenStudy (anonymous):

Ok, but I am very confused how to write this all out and solve it...

OpenStudy (mathstudent55):